OpenStudy (samigupta8):

Let f(x) be a differentiable function and G be the graph of f(x).Let P={a,f(a)} be a point on G closest to (0,0).Then f(a)f'(a)=?

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (irishboy123):

use the distance formula and minimise distance between a generalised point (x, f) [where f = f(x)] and the Origin by using \(s^2 = x^2 + f^2\) minimising \(\implies 2 s s' = 2x + 2f f' = 0\) at (a, f(a)) see it? if the line goes through the origin, you have a problem, i think, but not sure you need to worry about that

OpenStudy (samigupta8):

Actually i did this problem considering this. Slope of tangent at point A is f(a)/a (since it is close to origin so it must pass through origin). Now question says f(a)f'(a) so we can substitute it here. THAT BECOMES f(a)^2/a. Now , i assume that f(a) is a only. So got the ans to be a.

OpenStudy (irishboy123):

i think it's -a FWIW

OpenStudy (samigupta8):

Your ans is correct. Will you please tell how did get it?

OpenStudy (samigupta8):

You*

Parth (parthkohli):

it's simply a minimisation problem. find the distance in terms of \(x\), and then minimise the function. \[d(x) = \sqrt{x^2+y^2}=\sqrt{x^2 + f^2(x)}\]now we have to minimise the above expression. we can also say that we're minimising \(g(x) = x^2 + f^2(x)\). that will be true when \(g'(x) = 0\). now \(g'(x) = 2x + 2f(x)f'(x) = 0\). we're given that this equation holds at \(x = a\). this means\[2a + 2f(a)f'(a) = 0 \Rightarrow f(a)f'(a) = -a \]